Two rods A and B , each of equal length but different materials are suspended from a common support. The rods A and B can support a max load of W1 = 600 N and W2 = 6000 N, respectively. If their cross section areas are A1 = 10 mm 2 and A2 = 1000 mm2, respectively, then Identify the stronger material.

Grade:12

1 Answers

AskiitianExpert Pramod-IIT-R
14 years ago
Dear Student Mr. Sanchit; the stronger material is which one who can sustain higher stress value. this property is not depends on total load. So first of all , we have to calculate the maximum stress which material can sustain. stress = load/cross section area stress1 = 600/10 = 60 N/mm2 stress2 = 6000/1000 = 6N/mm2 so stress 1 is higher than stress 1 So materials 1 can sustain at higher stress. so stronger material is 1. Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE & AIEEE preparation. All the best. Regards, Askiitians Experts Pramod Kumar IITR Alumni